OpenStudy (anonymous):

1. What is the fourth term in the expansion of (x + 4y)^6? a) 7680x^3y^3 b) 3840x^2y^4 c)1440x^2y^4 d)1280x^3y^3 2. A bag contains 6 red marbles, 4 blue marbles, 7 green marbles, and 3 yellow marbles. A marble is drawn from the bag and is not replaced. Then a second marble is drawn. What is the probability that both marbles drawn were green? Option A: 49/380 Option B: 49/400 Option C: 21/190 Option D: 21/200 3. A bag contains 3 red marbles, 4 blue marbles, and 6 green marbles. Which pair of events are complementary events? a) choosing a red marble; choosing a blue marble b)choosing a re

OpenStudy (anonymous):

1) D

OpenStudy (anonymous):

as t_4=6C3(x)^3(4y)^3=1280x^3y^3

OpenStudy (anonymous):

2) 7/20*6/19=21/190 ans C

OpenStudy (anonymous):

yes easy

OpenStudy (anonymous):

hello chr34 have u understood??

OpenStudy (anonymous):

Do you know the last one?

OpenStudy (anonymous):

question is incomplete...

OpenStudy (anonymous):

yes i also saw that some thing was missing there

OpenStudy (anonymous):

Oops, #3 options: a)choosing a red marble; choosing a blue marble b) choosing a red marble; choosing a marble that is not red c) choosing a blue marble; choosing a marble that is not green d)choosing a green marble; choosing a red marble

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

b)

myininaya (myininaya):

dip can you tell me how you got the first one?